# Manga109Dialog: A Large-scale Dialogue Dataset for Comics Speaker Detection

### v1

@@ 4.4 Evaluation Metrics
The earliest and the most widely accepted evaluation metric of SGG models is Recall@K lu2016visual . To calculate this, we first need to compose the set of subject-predicate-object into a triplet. For each triplet, we compute the prediction score by multiplying each score. We sort the scores and select the top $K$ triplets. We then compute the triplet recalls. We consider the triplet a correct prediction when all three elements are labeled correctly.
In classic SGG tasks, $K$ is usually predefined as 20, 50, or 100. However, since the number of texts fluctuates significantly from image to image, using a fixed $K$ value does not provide a fair comparison. Take Figure 5 as an example. When a large $K$ is fixed, Recall@K can cover almost all major combinations of objects if the number of texts is small, which makes the accuracy very close to 100% (left side of Figure 5). When a small $K$ is fixed, the score is always low if there are many texts (right side in Figure 5).
To compare our proposed method with rule-based methods more appropriately, we set $K$= #text and introduce a new evaluation metric called Recall@(#text), which is the Recall of selected predictions that can cover all texts on the page. For each text, we choose the triplet containing it with the highest score as the prediction. For the cases where there exist $N$ speakers for a single text, we select the top $N$ triplets as predictions. We use Recall@(#text) to evaluate the experimental results in Section 5.3.
Figure 5: Comparison of using Recall@K and Recall@(#text). ©Arai Satoshi. ©Akamatsu Ken.
@@ 5.1 Experimental Setup
Baseline:
The simplest rule-based method is “shortest distance”, where we assume the text is spoken by the closest character. A slightly modified version is described as “frame distance”. We prioritize characters in the same frame as the text. If there are no characters in the frame, we make predictions using the same rule as “shortest distance”. In these two approaches, the amount of predictions is the same as that of texts. Therefore, there is no need to select the top $K$ predictions. Besides, we proposed a deep learning-based baseline, where the relationship score was calculated only by feeding the union feature $\textbf{f}_{i,j}$ obtained from Eq. 4 into a fully connected layer.
Datasets:
We divided Manga109Dialog into training and test sets with ratios of 70% and 30%, respectively. Furthermore, we divided the annotations into three levels of difficulty mentioned in Section 3.2. Besides, we tested our model on Abe’s dataset. Since the annotation in Abe’s dataset could not identify the particular bounding box of the speaker, we did an additional step and determined it by “frame distance”.
Evaluation:
In the task of SGG, the performance of the models can be evaluated in three protocols. (1) Predicate Classification (PredCls): predicting the relationship between two objects ($\{r_{i\to j}\}$), given an image ($I)$, object bounding boxes ($\{\mathbf{b}_{i}\}$), and object labels ($\{l_{i}\}$). (2) Scene Graph Classification (SGCls): predicting object labels and relationships, given an image and bounding boxes. (3) Scene Graph Detection (SGDet): detecting object bounding boxes and predicting their labels and relationships to generate an entire scene graph using only the given image. We executed the model for these three tasks to evaluate our proposed method.
5.2 Implementation Details
Because object detection is the most time-consuming step in speaker prediction, we pre-trained a Faster-RCNN model as our object detector. Following the previous studies in SGG tang2020unbiased ; suhail2021energy , we applied the Faster-RCNN with a ResNeXt-101-FPN backbone. We then froze the model as the object detector of our SGG models. Additionally, we pre-trained another Faster-RCNN as the frame detector. The Mean Average Precision (mAP) of the two models reached 86.09% and 96.48%, respectively. We present the details and results of our pre-trained object detector and frame detector in our supplementary material.
We trained our models on a single NVIDIA A100 GPU (60 GiB), with a batch size of 4. We used SGD as an optimizer and set the initial learning rate to $4\times 10^{-3}$. In the stage of SGG, we trained our SGG model on a single NVIDIA A100 GPU (60 GiB), with a batch size of 4. We optimized the model using SGD with an initial learning rate of $4\times 10^{-2}$. Our loss was the sum of the cross entropy for objects and relationships.
Figure 6: Recall@(#text) for PredCls.
@@ 5.3 Quantitative Results
We ran our model under three tasks (PredCls, SGCls, SGDet). Because our study focuses on speaker prediction, we only show the results under PredCls in Table 6, meaning that the model only needs to find speaker-to-text pairs. See the supplementary material for the results under SGCls and SGDet. From Table 6, we can observe a significant improvement in our approach compared to the rule-based approach, especially on Hard. That was because the speaker in Hard is not in the same frame as the text, while “frame distance” gives preference to the character in the same frame. Therefore, other characters are in the frame to which the text belongs, “frame distance” is 100% likely to be wrong. Besides, the model using reading order information performed better than that without it.
In addition to our dataset, we also tested our model on Abe’s dataset under the task of PredCls. The results can be find in our supplementary material.
@@ 5.4 Qualitative Results
To better understand the generated scene graphs visually, we visualized the predictions. Figure 7 shows Recall@(#text) under PredCls. The green lines represent correct predictions, and the red lines represent incorrect predictions. On Easy, our model outperformed the “frame distance” method regardless of whether frame reading orders were used. For texts in Easy, if the speaker was not the closest character to the text, rule-based methods are certain to make an incorrect prediction, as shown in the red lines. Moreover, the results on Hard show that the proposed method of introducing frame orders exhibited a high accuracy in prediction and was able to handle more challenging cases. We show some additional examples in the supplementary material.
@@ 5.5 Challenges and Future Work
Since speaker prediction is somewhat subjective, more annotators are needed to enhance the reliability and validity of Manga109Dialog and enable it to be used more effectively for research and development in the field of comics analysis.
Besides, although our proposed method demonstrated significant improvements compared to conventional methods, there were still some cases where predictions failed. Two situations were primarily challenging for the machine. The first case was that the speaker was not in the closest position to the text. Though our model can cope with this situation to some degree, it was not always able to make correct predictions. The other case was that texts spoken by different characters appear alternately. Because no visual information can be used for prediction, this may be relatively difficult even for humans. The straightforward way to handle these cases would be to introduce Natural Language Processing (NLP) models. However, this research aimed to show to what extent speaker detection can be achieved only with visual information. The results of the present study provide a baseline needed to incorporate NLP in future research.
@@ 6 Conclusion
(a) Predictions on Easy. ©Gasan, ©Saijo Shinji.
(b) Predictions on Hard. ©Saki Kaori.
Figure 7: Examples of predictions made by rule-based method and proposed method. The green lines represent correct predictions, while the red lines represent incorrect predictions.
In this study, we have constructed a large-scale dialogue dataset called Manga109Dialog based on the shortcomings of existing annotations. we have presented a novel approach that applies SGG models to comics speaker detection, thereby creating a benchmark for deep learning-based methods where none existed before. Due to comics’ unique features, we have introduced frame reading order to help predict the speaker. To properly evaluate the experimental results, we have further proposed a new evaluation metric and tested our model on different test sets. The results indicate the reliability of Manga109Dialog, as well as the superior performance of our deep learning-based approach over existing rule-based methods. These findings highlight the potential of scene-graph-based methods in this field of comics speaker detection. We believe our dataset and model will have broader applicability in the digital processing of comic. Our work lays the groundwork for further investigations into the effective combination of SGG models and other techniques such as NLP, offering new insights for future research in this area.

### v2

@@ 4.4 Evaluation Metrics
The earliest and the most widely accepted evaluation metric for SGG models is Recall@K, traditionally predefined at 20, 50, or 100 [19]. However, since the number of texts in comics fluctuates significantly from image to image, using a fixed $K$ value does not provide a fair comparison. Take Fig. 5 as an example. When a large $K$ is fixed, Recall@K can cover almost all major combinations of objects if the number of texts is small, which makes the accuracy very close to 100% (left side of Fig. 5). When a small $K$ is fixed, the score is always low if there are many texts (right side in Fig. 5).
To establish a more reliable benchmark for the task of comics speaker detection, we introduce a new evaluation metric called Recall@(#text), which measures the Recall of selected predictions covering all texts on a page. For each text, we choose the subject-predicate-object triplet containing it with the highest score as the prediction. In cases where a single text has $N$ speakers, we select the top $N$ triplets. We consider a triplet correct if the object regions are detected with an IoU greater than 0.65 and correctly labeled.
Fig. 5: Comparison of using Recall@K and Recall@(#text). Courtesy of Arai Satoshi, Akamatsu Ken.
@@ 5.1 Experimental Setup
Baseline: The simplest rule-based method is “shortest distance”, where we assume the text is spoken by the closest character. A slightly modified version is described as “frame distance”. We prioritize characters in the same frame as the text. If there are no characters in the frame, we make predictions using the same rule as “shortest distance”. Besides, we propose a deep learning-based baseline, where the relationship score was calculated only by feeding the union feature $\textbf{f}_{i,j}$ obtained from Eq. 4 into a fully connected layer.
Datasets: We divided Manga109Dialog into training and test sets with ratios of 70% and 30%, respectively. Furthermore, to see our model’s behavior on different styles of comics or languages, we tested it on eBDtheque, which comprises 100 images from French, American, and Japanese comics [20].
Evaluation: In the task of SGG, the performance of the models can be evaluated in three protocols. (1) Predicate Classification (PredCls): predicting the relationship between two objects, given an image $I$, object bounding boxes $\{\mathbf{b}_{i}\}$, and object labels $\{l_{i}\}$. (2) Scene Graph Classification (SGCls): predicting object labels and relationships, given an image and bounding boxes. (3) Scene Graph Detection (SGDet): detecting object bounding boxes and predicting their labels and relationships using only the given image. We executed the model for these three tasks to evaluate our proposed method.
Implementation Details: Because object detection is the most time-consuming step, we pre-trained a Faster-RCNN model as our object detector. Additionally, we pre-trained another Faster-RCNN as the frame detector. The Mean Average Precision (mAP) of the two detectors reached 86.09% and 96.48%, respectively. We present the details of them in the supplementary material. In the stage of SGG, we trained our model on a single NVIDIA A100 GPU, using the SGD optimizaer with a batch size of 4. We used the WarmupReduceLROnPlateau as the learning scheduler and set the initial learning rate to $4\times 10^{-2}$. The loss was the sum of the cross entropy for objects and relationships.
Table 2: Recall@(#text) for PredCls.
@@ 5.2 Quantitative Results
We ran our model under three tasks (PredCls, SGCls, SGDet). Because our study focuses on speaker prediction, we only show the results under PredCls in Table 2, meaning that the model only needs to find speaker-to-text pairs. See the supplementary material for the results under SGCls and SGDet.
Table 2 shows an improvement in our approach over rule-based approaches, especially on Hard. That was because the speaker in Hard is not in the same frame as the text, while “frame distance” gives preference to the character in the same frame. Therefore, other characters are in the frame to which the text belongs, “frame distance” is 100% likely to be wrong. Besides, the model using reading order information performed better than that without it.
Additionally, we directly tested our model on eBDtheque. Our model achieved 77.45% accuracy under PredCls, slightly below “frame distance”. This is mainly due to style variations in old European and American comics, where the speaker and text are often in the same frame. More details are available in the supplementary material.
@@ 5.3 Qualitative Results
To better understand the generated scene graphs, we visualized the predictions. Fig. 6 shows Recall@(#text) under PredCls. The green lines represent correct predictions, and the red lines represent incorrect predictions. For texts in Easy, if the speaker was not the closest character to the text, rule-based methods are certain to make an incorrect prediction, while our model outperformed the “frame distance” method regardless of whether frame reading orders were used. Moreover, the results on Hard show that the proposed method of introducing frame orders exhibited a high accuracy in prediction and was able to handle more challenging cases. We show some additional examples in the supplementary material.
(a) Predictions on Easy.
(b) Predictions on Hard.
Fig. 6: Examples of predictions made by rule-based method and proposed method. Courtesy of Gasan, Saijo Shinji, Saki Kaori. The green lines represent correct predictions, while the red lines represent incorrect predictions.
@@ 5.4 Challenges and Future Work
Although our proposed method demonstrated improvements compared to conventional methods, there were two situations primarily challenging. (1) The speaker was not in the closest position to the text. Though our model can cope with this situation to some degree, it did not always make correct predictions. (2) Texts spoken by different characters appear alternately. Because no visual information can be used for prediction, this may be relatively difficult even for humans.
The straightforward way to handle these cases would be to introduce Natural Language Processing (NLP) models. However, this research aimed to show to what extent speaker detection can be achieved only with visual information. The results of the present study provide a baseline needed to incorporate NLP in future research.
@@ 6 Conclusion
In this study, we developed Manga109Dialog, a large-scale dialogue dataset that significantly advances the field of comics analysis. We pioneered a novel approach by applying SGG models to the task of comics speaker detection, establishing a benchmark for deep learning-based methods where none existed before. The results not only demonstrate the reliability of Manga109Dialog but also show the superior efficacy of our scene-graph-based approach. Our work paves the way for future research, particularly in exploring the effective combination of SGG models and other techniques such as NLP, offering new insights in this field.
Appendix A Dataset Construction Details
This section shows the construction details of Manga109Dialog.
A.1 Annotation Rules
In this section, we use examples to show more details of annotation rules stated in Section 3.1.
First, let us introduce the basic rules. Although the images in Manga109 are double-page spreads, from the semantic point of view, we take every single page as a unit. If the number of the character who is saying the text is only one on the page, link the bounding box of the character to the text regardless of the position. If the number of the character who is saying the text is more than one on the page, the bounding box of the character in the same frame as the text is linked in priority. If the character is not in the same frame, it is determined by the reading order (from top to bottom, right to left). Take the rightmost text of Fig. A (a) as an example. Since there is no character in the frame with it, it is linked to the boy who appears in the middle frame according to the reading order.
(a) If the speaker is not in the same frame, it is determined by the reading order.
(b) If the texts are in the last frame of the page, they are linked to the character in the second-to-last frame.
Fig. A: Examples where the speaker is not in the same frame as the text. Courtesy of Inokuma Shinobu, Kurita Riku.
However, there exist exceptions. For texts in the last frame of the page, since we only consider the single page, we link the character in the second-to-last frame to the text instead of the character in the next page. One example is shown in Fig. A (b).
Next, let us consider some special cases. Take Fig. B (a) as an example. When there is more than one speaker for a single text, we link the text to the bounding boxes of all speakers.
Besides, for texts said by “Others”, if we can specify the bounding box of the speaker (such as Fig. B (b)), it is annotated; otherwise, it is excluded.
Although we make annotations on Manga109, not all texts are our annotation targets. For example, titles and descriptions are not annotated. For texts belonging to “Narration” or with unknown speakers, like those in Fig. C (a), we do not annotate them, either. Furthermore, when the speaker is not on the page, such as texts from phone conversations (Fig. C (b)) or letters, the text is not in the scope of annotation.
(a) An example of multiple speakers for a single text.
(b) An example of texts said by “Others”.
Fig. B: Some special cases. Courtesy of Akamatsu Ken, Kobayashi Yuki.
(a) Texts with unknown speakers are not our annotation targets.
(b) When the speaker is not on the page, the text is not annotated.
Fig. C: Texts not in the scope of annotation. Courtesy of Inokuma Shinobu, Akamatsu Ken.
A.2 Annotation Details
We outsourced the annotation of Manga109Dialog to a company with expert annotators, consisting of 8 operators and 3 inspectors. Operators used the annotation tool shown in Fig. D to link the speaker to the text. Then inspectors checked the annotations for omissions and mistakes. Creating the annotations took approximately three months, and work hours (except for inspections) exceeded 627 hours. The annotations for each comic are saved in an xml file, an example is shown in Fig. E.
A.3 Visualization Tool
After the annotation company completed the task, we confirmed and corrected their annotations. To check the annotations in a visual way, we designed a visualization tool. The interface is shown in Fig. F.
When pressing the “relationship” button, the visualization tool can show the regions of characters and texts with different colored rectangular boxes. Additionally, it can show characters’ names and speaker-to-text pairs. As shown in Fig. F, each orange line connects the center coordinates of the speaker and the text. When a text has $n$ speakers, there will be $n$ lines.
Fig. D: The interface of the annotation tool. Courtesy of Yoshi Masako.
Fig. E: Format of the saved annotations.
Fig. F: The interface of our visualization tool. Courtesy of Kato Masaki.
Appendix B Experimental Preparation
This section visually explains the details of our pre-trained object detector, frame detector, and frame order estimator.
(a) Ground truth.
(b) Detection results.
Fig. G: An example of detection results. Courtesy of Matsuda Naomasa.
B.1 Object Detector
We pre-trained our object detector to save run time of speaker prediction. Following the previous studies in SGG [12], we applied the Faster-RCNN with a ResNeXt-101-FPN backbone. We then froze the model as the object detector. We trained our models on a single NVIDIA A100 GPU, with a batch size of 4. We used SGD as an optimizer and set the initial learning rate to $4\times 10^{-3}$.
We set object classes to three: character, text and background. The mAP of our detector was 86.09%. The high accuracy may be related to the small number of object categories in our experiments. An example of detection results is shown in Fig. G. The left side of Fig. G shows the visualization of ground truth, and the right side shows the detected bounding boxes and predicted labels. We can also see from this example that our model has a high accuracy.
B.2 Frame Detector and Frame Order Estimator
We pre-trained another Faster-RCNN model as the frame detector, which was trained on a single NVIDIA A100 GPU, with a batch size of 4. We used SGD as an optimizer and set the initial learning rate to $4\times 10^{-3}$. We set object classes to two: frame and background and our frame detector achieved an mAP of 96.48%.
For detected frames, we used the algorithm described in Kovanen and Ikuta’s study [17, 18] to presume their reading orders. Given a set of frame bounding boxes, we first decide whether they are horizontally divisible into two parts. If so, we split them and repeat the first step for each set until we cannot horizontally split them. Then, we decide whether each set is vertically divisible into two parts. If so, we do so and repeat the first step. When the set of frames cannot be divided horizontally and vertically, we number the frames according to reading order (from top to bottom, from right to left).
An example of the estimation results is shown in Fig. H. When there is a high degree of overlap in a set of frames, we consider the set inseparable. All frames in this set are given the same order and are shown in orange dotted-line bounding boxes. For the example in Fig. H, the first frame covers half of the page in ground truth. Since the following three frames are inside the first frame, the first four frames are considered inseparable and share the same reading order. Therefore, although the detected frame is different from those in ground truth, the reading order is presumed correctly, so there is no negative impact on our experiments of speaker prediction.
(a) Ground truth.
(b) Detection results.
Fig. H: Results of frame detection and reading order estimation. Courtesy of Hikochi Sakuya.
Appendix C Rule-Based Baseline
Here, we recap simple rule-based approaches using our notations in Section 4.2.
The simplest baseline is “shortest distance”. That is, the score $r_{i\to j}$ is defined as follows.
$r_{i\to j}=\begin{cases}1&\mathrm{if}~{}\mathbf{b}_{i}~{}\mathrm{is~{}closest~{}to}~{}\mathbf{b}_{j},\\ &l_{i}=\mathtt{character},~{}~{}\mathrm{and}~{}~{}l_{j}=\mathtt{text}.\\ 0&\mathrm{else}.\end{cases}$ (a)
Here, we assume that $\{\mathbf{b}_{i}\}_{i=1}^{N}$ is given by $f_{\mathrm{detect}}$. This equation means that the text is spoken by the closest character. This is simple and can be computed rapidly, but cannot handle complex cases.
A slightly modified version is described as “frame distance”, which is defined as given below.
[…]
Here, we prioritize characters in the same frame as the text. If there are no characters in the frame, we make predictions using the same rule as “shortest distance”. In these two approaches, the amount of $r$ is the same as that of $\mathtt{text}$. Therefore, there is no need to select the top $K$ predictions.
Appendix D Ablation Studies
To evaluate our proposed method, we conducted ablation studies and ran our model under the following four patterns.
  • Rule-based: methods discribed in Section C. In “frame distance”, the frames are obtained from our frame detector.
  • Only w/ union feature: a deep learning-based simple baseline that calculates relationship scores only using the union feature $\textbf{f}_{i,j}$ obtained in Eq. 4, without using the fusion module $f_{\mathrm{fuse}}$ in Eq. 2.
  • Proposed w/o frame: predict speakers without frame information. In this case, the process is the same as the original SGG process.
  • Proposed w/ detected frame: use the frame bounding boxes detected by Faster-RCNN. This scenario is a practical use case. We detect both the frame positions and the reading orders.
  • Proposed w/ given frame: use the frame bounding boxes from Manga109’s annotations. This means that ground truth bounding boxes of frames are available. We then estimate the frame orders.
Recall@(#text) under three SGG tasks is shown in Table BBD.
From these tables, we can observe a significant improvement for the model with frame information over the model without it. Moreover, the model with given frames performs better than the model with detected frames overall, but “Proposed w/ detected frame” overperforms “Proposed w/ given frame” on Hard. It may be because the speaker and the text are not in the same frame in data on Hard, while our frame detector divided frames into smaller ones, making the estimated reading order more accurate.
Appendix E Experiments with Different Styles of Comics
To see the behavior of our model on different styles of comics or languages, we tested our model on eBDtheque [20], which comprises 100 images from French, American, and Japanese comics. Results under PredCls is shown in Table D. From the table, it may be observed that although our model was trained on Manga109, it also performed well on eBDtheque. Our method outperformed “shortest distance” but did not perform as well as “frame distance”. This was mainly because, in contrast to Japanese manga, the speaker and the text are almost always in the same frame in relatively old European and American comics. This experiment highlights the variations in comic styles across different countries and outlines directions for future research.
Table A: Recall@(#text) for PredCls.
Manga109Dialog
Method Easy Hard Total
Rule-based
      Short distance 71.43 22.72 63.41
      Frame distance 81.55 22.06 71.53
Deep learning-based
      Only w/ union feature 83.20 28.50 73.99
      Proposed w/o frame 84.73 29.55 75.48
      Proposed w/ detected frame 84.77 30.65 75.69
      Proposed w/ given frame 85.24 29.50 75.89
Table B: Recall@(#text) for SGCls.
Manga109Dialog
Method Easy Hard Total
Rule-based
      Short distance 71.06 23.71 63.09
      Frame distance 80.81 20.61 70.68
Deep learning-based
      Only w/ union feature 82.74 28.41 73.59
      Proposed w/o frame 84.42 29.12 74.92
      Proposed w/ detected frame 84.51 29.77 75.32
      Proposed w/ given frame 84.97 29.23 75.62
Table C: Recall@(#text) for SGDet.
Manga109Dialog
Method Easy Hard Total
Rule-based
      Short distance 53.11 14.85 46.67
      Frame distance 64.49 13.67 55.93
Deep learning-based
      Only w/ union feature 67.48 21.31 59.71
      Proposed w/o frame 67.84 21.43 60.04
      Proposed w/ detected frame 67.92 24.40 60.54
      Proposed w/ given frame 68.21 22.96 60.60
Table D: Experimental results on eBDtheque.
Method eBDtheque
Rule-based
      Short distance 77.34
      Frame distance 80.27
Deep learning-based
      Proposed (SGG) w/o frame 75.20
      Proposed (SGG) w/ frame 77.45
Appendix F Experiments with Other SGG Models
This section explains why we chose MOTIFS as the SGG model for our proposed pipeline. In addition to MOTIFS, we ran experiments using other SGG models. We changed the LSTMs in the fusion module $f_{\mathrm{fuse}}$ to TreeLSTMs of VCTree [21, 10] and Transformer [22, 12], respectively. Experimental results are shown in Table E.
From the table, we can see that the methods using the SGG models all perform better than the rule-based method and DL simple baseline. However, there is not a large gap between the results of the different SGG models. It is probably because the relationship is only one class in our experiments, and it is difficult for models such as LSTMs to exert their advantages.
Table E: Results using different SGG models.
Manga109Dialog
Method Easy Hard Total
Frame distance 81.55 22.06 71.53
DL simple baseline 83.20 28.50 73.99
Proposed SGG
      MOTIFS 84.73 29.55 75.48
      VCTree 83.30 28.85 74.16
      Transformer 83.76 28.63 74.51
Appendix G Qualitative Results
In this section, we show more examples of prediction results. Speaker predictions on Easy, Hard and Total are given in Fig. I, Fig. J, Fig. K and Fig. L, which show Recall@(#text) in the task of PredCls. The green lines represent correct predictions, while the red lines represent incorrect predictions.
As we can see, our proposed model shows better performance and can deal with more complex cases, especially when there is at least one character in the same frame as the text. In this case, if the speaker is not the closest character to the text (like Fig. I and the lower part in Fig. K) or the speaker is not in the same frame (like the upper part in Fig. K), “frame distance” will surely make a wrong prediction, while our model can ease this problem to some degree.
Fig. I: Predictions on Easy. Courtesy of Yabuno Tenya, Watanabe Tatsuya, Saijo Shinji. The green lines represent correct predictions, while the red lines represent incorrect predictions.
Fig. J: Predictions on Hard. Courtesy of Yabuno Tenya, Watanabe Tatsuya.
Fig. K: Predictions on Total. Courtesy of Tashiro Kimu.
(a) Frame distance.
(b) Proposed SGG w/o frame.
(c) Proposed SGG w/ frame.
Fig. L: Predictions on Total. Courtesy of Kurita Riku.